Suppose u, v is the element of V and || u || = || v || = 1 with < u,v > = 1: Prove

that u = v.


1
Expert's answer
2021-07-06T08:55:06-0400

It follows that "\cos \alpha=\frac{<u,v>}{||u||\cdot ||v||}=\frac{1}{1}=1," and hence "\alpha =0." Since the vectors "u" and "v" have the same direction and the same magnitude, we conclude that "u=v."